# High-Level Synthetic Data Generation with Data Set Archetypes

### v3

@@ 4.2 Minimax Classification Error Captures Clustering Difficulty
In Section 3.2, we defined the overlap between two clusters in terms of the error rate of the best minimax linear classifier. We verify that this notion of overlap conveys clustering difficulty by measuring clustering performance on data sets with different degrees of overlap. For this simulation, we consider data sets with two clusters drawn from an archetype we described as “two clusters with very different shapes in $p$D” [^5] This verbal description yields an archetype with parameters
{ ‘name’: ‘two_very_different_shapes_$p$d’, ‘n_clusters’: 2, ‘dim’: $p$, ‘n_samples’: 200, ‘aspect_ref’: 1.5, ‘aspect_maxmin’: 3.0, ‘radius_maxmin’: 3, ‘imbalance_ratio’: 2, ‘max_overlap’: 0.05, ’min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’] }
, where the dimensionality $p$ ranges across $[2,10,30,100]$. We vary max_overlap from $10^{-7}$ to 0.5, while setting min_overlap = max_overlap/10. For each overlap setting, we generate 100 distinct data sets and evaluate the average clustering performance of hierarchical clustering, quantified in terms of adjusted mutual information (AMI) and adjusted Rand index (ARI), as in the benchmark of Section 4.1. We choose hierarchical clustering because it is computationally efficient and performed well in the benchmark. We repeat this process twice, where in the second run we make clusters non-convex by applying the distort function.
Figure 9 confirms that clustering difficulty rises with increasing overlap. Figure 10 shows the same in the case of non-convex clusters, suggesting that applying distort maintains the desired relationship between overlap and clustering difficulty. Additionally, both figures show how our cluster overlap relates to the silhouette score, a popular metric for quantifying clustering difficulty (Rousseeuw, (1987), Shand et al., (2019)). At a fixed value of max_overlap, the silhouette score decreases markedly with a rise in dimensionality. This is not an artifact of our overlap measure, since plotting clustering performance vs silhouette score shows a similar dependence on dimensionality (not shown). This makes sense since the silhouette score is based on the difference of Euclidean distances, and distances between points tend to become more similar in high dimensions (Aggarwal et al., (2001)).
@@ 5 Related Work
The popular scikit-learn library for machine learning (Pedregosa et al., (2011)) offers several functions for creating synthetic clusters. Among these, some are aimed at reproducing canonical 2D toy data sets like concentric circles and nested moons (make_moons, make_circles), while others focus on sampling multivariate normal clusters. These functions offer some valuable features, such as the ability to create datasets with informative and redundant features, as in the make_classification function. However, they do not control overall geometric characteristics of the data.
The data mining framework ELKI (Schubert and Zimek, (2019)) provides a synthetic data generator based on specifying probabilistic models in XML files. This XML-based approach makes it easy to reproduce benchmarks by sharing the underlying XML files. Drawing inspiration from this work, we have implemented an Archetype.describe function allowing users to easily share collections of data set archetypes as JSONL files.
The generators OCLUS (Steinley and Henson, (2005)) and GenRandomClust (Qiu and Joe, (2006)) focus on providing more sophisticated overlap control compared to previous generators. GenRandomClust extends the generator of Milligan and Cooper, (1985) by managing overlaps between clusters with different ellipsoidal shapes and arbitrary spatial orientations. Similar to our classification error-based notion of cluster overlap, their method finds an optimal separation direction between two clusters. To enforce the desired degree of overlap, the algorithm initially places cluster centers on a scaffold, then scales individual edges of the scaffold up or down to meet the overlap constraint. The method supports making all cluster shapes more or less elongated, but does not otherwise provide high-level control over data set geometry. Moreover, the scaling operations undertaken to manage cluster overlaps implicitly sacrifice control over cluster volumes (which, in repliclust, can be managed independently).
OCLUS (Steinley and Henson, (2005)) quantifies cluster overlap in terms of the shared density between two clusters. The generator uses analytical formulas for integrals of several interesting probability distributions (including exponential, gamma, and chi-square), thereby effectively managing overlaps between non-normal clusters. As a result, the method is limited to treating all dimensions independently, so that cluster distributions simplify as the products of marginals. The paper contains a valuable discussion of the distinction between marginal and joint cluster overlaps (of which their software supports both).
@@
MDCGen (Iglesias et al., (2019) is a feature-rich generator that supports many desiderata in cluster analysis, such as overlap control, different probability distributions, subspace clusters, and the ability to add noise points. In particular, it is nice to be able to place noise points away from the clusters, which is made possible by the grid-based strategy for placing cluster centers. MDCGen does not target the overall geometric characteristics of synthetic data sets, instead giving users low-level control enabling extensive configurability. For example, managing the overlap between clusters involves setting compactness coefficients, grid granularity, and overall scale, compared to only tweaking max_overlap in repliclust.666max_overlap may also have to be tweaked but it can usually stay at max_overlap/10 or a similar value. In the words of the authors, “to enable covering a broad range of dataset possibilities, the parameters are multiple and some training for tuning the tool is required.”
Finally, the HAWKS generator (Shand et al., (2019)) controls cluster overlaps using an evolutionary algorithm that evolves the means and covariance matrices of multivariate normal distributions. The paper applies this framework to create data sets with a user-specified silhouette score representing clustering difficulty (Rousseeuw, (1987)). In principle, the evolutionary framework can be extended to attain desired high-level geometric characteristics. Of these, the authors consider two examples, cluster overlaps and elongations (the latter relating to our notion of cluster aspect ratio, as listed in Table 1). An interesting aspect of HAWKS is the ability to generate data sets that maximize the performance difference between two clustering algorithms. This feature is especially useful in two dimensions, since we can then visually examine the data sets to better understand when each algorithm succeeds or fails.
@@ Acknowledgements
MJZ would like to acknowledge support from Matt Thomson and members of the Thomson Lab at Caltech. In addition, MJZ thanks Dante Roy Calapate for providing his computer monitor during the early stages of this project.
@@
We prove Theorem 1 of Section 3.2. Additionally, we provide an analogous result for the cruder “center-to-center” approximation of cluster overlap.
@@
The default architecture for the neural network used in the distort function of Section 3.4 is a feed-forward network consisting (in order) of a linear embedding, 16 repeated feed-forward blocks (each consisting of a fully connected linear layer followed by layer normalization and Tanh activation), and a linear projection. Importantly, we tie the weights of the embedding and projection layers, so that the projection weights are the transpose of the embedding weights (Inan et al., (2017)). The default hidden dimensionality is 128, so that the embedding layer maps a data set archetype’s dimension to 128. The internal feed-forward blocks preserve this hidden dimension, and the final projection layer maps it back to the archetype’s dimension.
@@ 1. Prompt template mapping archetype description to high-level geometric parameters:
distributions: list[str], determines the probability distributions to use for the clusters; the available distributions are normal’, standard_t’, exponential’, beta’, uniform’, chisquare’, gumbel’, weibull’, gamma’, pareto’, f’, and lognormal
@@
Archetype JSON: {"distributions": [’normal’, standard_t’, exponential’, beta’, uniform’, chisquare’, gumbel’, weibull’, gamma’, f’, and lognormal’]}
@@
Archetype JSON: {"distributions": [’normal’, exponential’, beta’, uniform’, chisquare’, gumbel’, weibull’, gamma’, f’, and lognormal’]}
@@
"distributions": [’normal’, exponential’, beta’, uniform’, chisquare’, gumbel’, weibull’, gamma’, f’, and lognormal’]}

### v4

@@ 4.2 Minimax Classification Error Captures Clustering Difficulty
In Section 3.2, we defined the overlap between two clusters in terms of the error rate of the best minimax linear classifier. We verify that this notion of overlap conveys clustering difficulty by measuring clustering performance on data sets with different degrees of overlap. For this simulation, we consider data sets with two clusters drawn from an archetype we described as “two clusters with very different shapes in $p$D”. This verbal description yields an archetype with parameters
{ name’: ‘two_very_different_shapes_$p$d’, n_clusters’: 2, ‘dim’: $p$, n_samples’: 200, ‘aspect_ref’: 1.5, aspect_maxmin’: 3.0, ‘radius_maxmin’: 3, imbalance_ratio’: 2, ‘max_overlap’: 0.05, ’min_overlap’: 0.001, distributions’: [‘normal’, exponential’] },’
where the dimensionality $p$ ranges across $[2,10,30,100]$. We vary max_overlap from $10^{-7}$ to 0.5, while setting min_overlap = max_overlap/10. For each overlap setting, we generate 100 distinct data sets and evaluate the average clustering performance of hierarchical clustering, quantified in terms of adjusted mutual information (AMI) and adjusted Rand index (ARI), as in the benchmark of Section 4.1. We choose hierarchical clustering because it is computationally efficient and performed well in the benchmark. We repeat this process twice, where in the second run we make clusters non-convex by applying the distort function.
Figure 9 confirms that clustering difficulty rises with increasing overlap. Figure 10 shows the same in the case of non-convex clusters, suggesting that applying distort maintains the desired relationship between overlap and clustering difficulty. Additionally, both figures show how our cluster overlap relates to the silhouette score, a popular metric for quantifying clustering difficulty (Rousseeuw,, 1987; Shand et al.,, 2019). At a fixed value of max_overlap, the silhouette score decreases markedly with a rise in dimensionality. This is not an artifact of our overlap measure, since plotting clustering performance vs silhouette score shows a similar dependence on dimensionality (not shown). This makes sense since the silhouette score is based on the difference of Euclidean distances, and distances between points tend to become more similar in high dimensions (Aggarwal et al.,, 2001).
@@ 5 Related Work
The popular scikit-learn library for machine learning (Pedregosa et al.,, 2011) offers several functions for creating synthetic clusters. Among these, some are aimed at reproducing canonical 2D toy data sets like concentric circles and nested moons (make_moons, make_circles), while others focus on sampling multivariate normal clusters. These functions offer some valuable features, such as the ability to create datasets with informative and redundant features, as in the make_classification function. However, they do not control overall geometric characteristics of the data.
The data mining framework ELKI (Schubert and Zimek,, 2019) provides a synthetic data generator based on specifying probabilistic models in XML files. This XML-based approach makes it easy to reproduce benchmarks by sharing the underlying XML files. Drawing inspiration from this work, we have implemented an Archetype.describe function allowing users to easily share collections of data set archetypes as JSONL files.
The generators OCLUS (Steinley and Henson,, 2005) and GenRandomClust (Qiu and Joe,, 2006) focus on providing more sophisticated overlap control compared to previous generators. GenRandomClust extends the generator of Milligan and Cooper,, 1985 by managing overlaps between clusters with different ellipsoidal shapes and arbitrary spatial orientations. Similar to our classification error-based notion of cluster overlap, their method finds an optimal separation direction between two clusters. To enforce the desired degree of overlap, the algorithm initially places cluster centers on a scaffold, then scales individual edges of the scaffold up or down to meet the overlap constraint. The method supports making all cluster shapes more or less elongated, but does not otherwise provide high-level control over data set geometry. Moreover, the scaling operations undertaken to manage cluster overlaps implicitly sacrifice control over cluster volumes (which, in repliclust, can be managed independently).
OCLUS (Steinley and Henson,, 2005) quantifies cluster overlap in terms of the shared density between two clusters. The generator uses analytical formulas for integrals of several interesting probability distributions (including exponential, gamma, and chi-square), thereby effectively managing overlaps between non-normal clusters. As a result, the method is limited to treating all dimensions independently, so that cluster distributions simplify as the products of marginals. The paper contains a valuable discussion of the distinction between marginal and joint cluster overlaps (of which their software supports both).
@@
MDCGen (Iglesias et al.,, 2019 is a feature-rich generator that supports many desiderata in cluster analysis, such as overlap control, different probability distributions, subspace clusters, and the ability to add noise points. In particular, it is nice to be able to place noise points away from the clusters, which is made possible by the grid-based strategy for placing cluster centers. MDCGen does not target the overall geometric characteristics of synthetic data sets, instead giving users low-level control enabling extensive configurability. For example, managing the overlap between clusters involves setting compactness coefficients, grid granularity, and overall scale, compared to only tweaking max_overlap in repliclust (min_overlap may also have to be tweaked but it can usually stay at max_overlap/10 or a similar value). In the words of the authors, “to enable covering a broad range of dataset possibilities, the parameters are multiple and some training for tuning the tool is required.”
Finally, the HAWKS generator (Shand et al.,, 2019) controls cluster overlaps using an evolutionary algorithm that evolves the means and covariance matrices of multivariate normal distributions. The paper applies this framework to create data sets with a user-specified silhouette score representing clustering difficulty (Rousseeuw,, 1987). In principle, the evolutionary framework can be extended to attain desired high-level geometric characteristics. Of these, the authors consider two examples, cluster overlaps and elongations (the latter relating to our notion of cluster aspect ratio, as listed in Table 1). An interesting aspect of HAWKS is the ability to generate data sets that maximize the performance difference between two clustering algorithms. This feature is especially useful in two dimensions, since we can then visually examine the data sets to better understand when each algorithm succeeds or fails.
@@ Acknowledgements
We thank the anonymous referees for their feedback, which has significantly improved the paper. In addition, MJZ would like to acknowledge support from Matt Thomson and members of the Thomson Lab at Caltech. Finally, MJZ thanks Dante Roy Calapate for providing his computer monitor during the early stages of this project.
Compliance with Ethical Standards
The project did not involve any studies with human or animal participants.
Conflicts of Interest
The authors report no conflicts of interest.
Funding
This project is not tied to a specific grant.
Data Availability
Data and code employed in writing the paper is available on GitHub in the repository github.com/mzelling/repliclust-analysis.
@@
We prove Theorem 1 of Section 3.2. Additionally, we provide an analogous result for the simpler “center-to-center” approximation of cluster overlap.
@@
The default architecture for the neural network used in the distort function of Section 3.4 is a feed-forward network consisting (in order) of a linear embedding, 16 repeated feed-forward blocks (each consisting of a fully connected linear layer followed by layer normalization and Tanh activation), and a linear projection. Importantly, we tie the weights of the embedding and projection layers, so that the projection weights are the transpose of the embedding weights (Inan et al.,, 2017). The default hidden dimensionality is 128, so that the embedding layer maps a data set archetype’s dimension to 128. The internal feed-forward blocks preserve this hidden dimension, and the final projection layer maps it back to the archetype’s dimension.
@@ 1. Prompt template mapping archetype description to high-level geometric parameters:
distributions: list[str], determines the probability distributions to use for the clusters; the available distributions are ’normal’, ’standard_t’, ’exponential’, ’beta’, ’uniform’, ’chisquare’, ’gumbel’, ’weibull’, ’gamma’, ’pareto’, ’f’, and ’lognormal’
@@
Archetype JSON: {"distributions": [’normal’, ’standard_t’, ’exponential’, ’beta’, ’uniform’, ’chisquare’, ’gumbel’, ’weibull’, ’gamma’, ’f’, and ’lognormal’]}
@@
Archetype JSON: {"distributions": [’normal’, ’exponential’, ’beta’, ’uniform’, ’chisquare’, ’gumbel’, ’weibull’, ’gamma’, ’f’, and ’lognormal’]}
@@
"distributions": [’normal’, ’exponential’, ’beta’, ’uniform’, ’chisquare’, ’gumbel’, ’weibull’, ’gamma’, ’f’, and ’lognormal’]}